OpenStudy (anonymous):

@amistre64 How would you put (n-4)(n-3)(n-2) into factorial form, i am soo confusedddd

OpenStudy (amistre64):

do you know what factorial form is? what a factorial is defined to look lke?

OpenStudy (anonymous):

yes, it has a "!" at the end of a term

OpenStudy (amistre64):

well yeah, but rather, how do we define n! ??

OpenStudy (anonymous):

n!=n(n-1)(n-2)(n-3)...3*2*1

OpenStudy (amistre64):

lets use that then, since we have n-2, n-3, and n-4

OpenStudy (amistre64):

if we start with (n-2)! and divide it by some (n-r)! what would r have to be?

OpenStudy (anonymous):

uh, is it 5

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

\[\frac{(n-2)!}{(n-5)!}=(n-2)(n-3)(n-4)\]

OpenStudy (anonymous):

thanks :D

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

actually i have one more question

OpenStudy (anonymous):

In how many ways can he flag that has 3 sections be coloured if you have nine colours to choose from? Adjacent regions cannot have th

OpenStudy (anonymous):

I did 9P3, but im not sure if its right